# pH-Driven Fluorescence in Solid-State Lysine

### v1

@@ Abstract
Our experimental characterization clearly shows that both pH and counterion modulate the emission of solid‑state Lys aggregates, highlighting the influence of protonation and aggregation on fluorescence. However, these techniques cannot directly resolve the molecular origins of this effect, specifically, how protonation state shapes the electronic structure and ES dynamics. To gain molecular insight into this, we now move to discussing the theoretical results that emerge from our non-adiabatic molecular dynamics simulations performed on three Lys crystal models corresponding to distinct protonation states, as shown in Fig. 4.
@@
Structural differences among the lysine crystals reflect protonation-dependent variations that mirror the pH effects observed in aggregate morphologies, supporting their use as proxies to study solid-state Lys photophysics. As we will see shortly, these models also correctly reproduce the trends in the emission observed as a function of pH. L-Lys sulfate (LLS, Fig. 4a) adopts a folded conformation stabilized by a dense HB network between SO and the -NH and CO2H groups of lysine36, 47, with HBs exclusively between Lys and sulfate and no direct Lys-Lys interactions due to electrostatic repulsion. In L-Lys monohydrochloride dihydrate (LLMHCl, Fig. 4b), all amines are protonated, the C-terminus is deprotonated, and both Lys–Lys and Lys–water HBs are observed. Cl- counterions further stabilize the structure via electrostatic interactions with both the - and -amino groups37, 48. The zwitterionic L-Lys hemihydrate (LLH, Fig. 4c) forms layers of head-to-head hydrogen bonded dimers through -NHCO interactions, a network of interactions previously linked to intrinsic fluorescence38, 49, 4.
@@
For each crystal, we performed 500 fs NAMD simulations, within a QM/MM framework, on 50 statistically independent initial conditions (see the Computational Details section in the SI for details). In Fig. 4a–c, the lysine dimers (rendered with glossy CPK spheres) represent the QM regions used in our simulations. After 500 fs, most trajectories remained in the first ES for LLH (pH = 10) and LLMHCl (pH = 7), namely 98 and 92, respectively, whereas only 28 remained in S1 for LLS (pH = 1), as shown in Fig. 4d (orange plot). This severe quenching at low pH mirrors the experimental EQE trend observed in Fig. 3f, reproduced here for direct comparison, highlighting the key role of protonation in modulating the ES lifetimes.
@@
In Fig. 4e, we report the fluorescence spectra of LLS (red line), LLMHCl (green line), and LLH (cyan line) averaged over the trajectories remaining in S1. Vertical lines indicate the averaged lowest absorption (excitation) transitions: blue (LLS), magenta (LLMHCl), and yellow (LLH). For all systems, excitation wavelengths cluster near nm, exhibiting a 120 nm ( eV) blue-shift relative to experiment. This discrepancy likely arises from a combination of known TDDFT limitations, such as the tendency of CAM-B3LYP to overestimate vertical excitation energies50, as well as neglect of nuclear quantum effects (NQEs), and structural defects or disorder51, 52. Several previous studies have shown that the inclusion of NQEs can alter both the peak absorption in organic systems by up to 1 eV, as well as affect the line-shape53, 54, 2, 55.
@@
Across protonation states, the transition consistently occurs from the highest occupied molecular orbital (HOMO) to the lowest unoccupied molecular orbital (LUMO), which exhibit n and character, respectively. An example for LLS is shown in Fig. 4f: the n-type HOMO localizes on the C-terminal carboxyl and partly on SO oxygen atoms, while the -type LUMO is on the same Lys monomer, primarily on the C-terminal oxygen and along the CCα bond. Similar localizations occur for LLMHCl (Fig. S6a) and LLH (Fig. S7a), with the n-type HOMO on CO oxygens and the -type LUMO on this group and adjacent Cα. Cl- ions show no active role in LLMHCl. The electron density alternates between the two QM Lys monomers, consistent with dimer symmetry, and localizes within H-bonding regions, linking local non-covalent interactions with photoexcited electronic properties.
@@
While the excitation energies show little pH dependence, emission spectra reveal more pronounced differences. The computed emission trend in Fig. 4e mirrors the one observed experimentally for Lys samples acidified with H2SO4 (Fig. 3f). The crystal representing pH 1 (LLS) shows the most blue-shifted emission ( nm), followed by LLMHCl (pH 7) and LLH (pH 10), both exhibiting emission maxima near 356 nm. Notably, the emission spectrum of LLMHCl displays an additional broad feature between 460–480 nm, originating from structures in which excited state distortions–particularly CCO2 deplanarization–are partially activated but insufficient to reach the conical intersection (CoIn). This mode, typically associated with non-radiative decay6, in this case gives rise to red-shifted emission from distorted yet emissive geometries. LLS emission instead exhibits a shoulder near the excitation energy, suggesting an ES minimum near the Franck-Condon region. Although emission energies remain overall blue-shifted by 50–80 nm (0.5-0.7 eV) relative to experiment, this discrepancy is significantly reduced compared to excitation, and spectral tails at longer wavelengths ( nm) are well captured for LLMHCl and LLH. In all three systems, the electronic transition retains the same n- character and orbitals localization as in the GS excitation (see Figs. S6-S8 in the SI).
Since the calculated emission spectra in Fig. 4e are normalized, this does not directly allow for a comparison of the intensities. We therefore investigated the oscillator strength (f) distributions for further insight. Normalized histograms and corresponding Gaussian fits for LLS, LLMHCl, and LLH are shown in Fig. S9a-c, with an overlay of the fitted curves in Fig. S9d. For the more fluorescent LLH (cyan) and LLMHCl (green), distributions peak below 0.001, with minor tails up to 0.005. These values indicate modest fluorescence intensity, well below that of typical aromatic fluorophores56, yet sufficient for detection (Fig. 4d). In contrast, LLS (red) shows a broader distribution with oscillator strengths reaching up to 0.01; however, the fluorescence signal is minimal due to the predominance of non-radiative decay pathways.
@@
Non-radiative decay on the excited state essentially involves fluctuations or distortions along vibrational modes that lead to the CoIn and subsequently, decay to the ground state. Previous studies2, 57, 3, 18, 6 have shown that proton transfer (PT), carbonyl stretching, and deplanarization of the amide group act as critical modes in modulating the ES lifetimes of peptide chains, amyloid-like peptides, and glutamine crystals. Constraining these modes raises the energy barrier to the conical intersection, prolonging the ES lifetime. More recently, Banerjee and co-workers7 linked the weaker fluorescence of cysteine crystals grown in H2O, as opposed to D2O, to activation of CαN and SH stretching modes. Guided by the n character of LLS (Fig. 4f), we first focused on vibrational modes involving the C-terminus, as shown in Figs. S10-S11 for the two QM Lys monomers. We then investigated the coordinates associated with the various HBs present in the system (Fig. S12 in the SI). From this combined set of vibrational modes, a small number of coordinates, shown in Fig. 5, emerge as the dominant funnels to the CoIn.
@@
The collective motion driving non-radiative decay in LLS (Fig. 5a) can be decomposed into two main contributions: (i) stretching of the CO bond coupled with deplanarization of the CαCO2 group, and (ii) PT coordinates involving the NH or OH groups of Lys and the sulfate oxygen. As shown in Fig. 5b, the CO bond elongates in S1 (orange distribution), increases further at the CoIn (dashed line), and contracts upon relaxation to S0 (blue distribution). A similar trend is observed for the second CO bond of the carboxylate group (Fig. S11c). This distortion is concerted with deplanarization of the amide group in S1, followed by replanarization after decay (Fig. S11d). These features are consistent with previously reported ES dynamics in non-emissive amino acid crystals, such as L-glutamine6.
@@
However, unlike glutamine–where CO stretching is triggered by transient strengthening (PT coordinate ) of a -NHCO hydrogen bond between monomers–LLS exhibits HBs exclusively with sulfate counterions. As a result, CO bond activation and proton transfer do not necessarily occur within the same structural unit. The primary HB-related mode promoting non-radiative decay involves transient weakening of an -NHSO HB in S1, particularly near the CoIn, with PT typically occurring upon relaxation to S0 (Figs. 5c and S12f). A secondary PT mode, shown in Fig. 5d, involves the C-terminal OH group and displays a slight transient HB strengthening in S1, with occasional proton sharing or transfer to the sulfate oxygen at the CoIn, and reversion after decay. A significant number of trajectories that reach the CoIn show PT coordinates below Å, consistent with a weakening of the interaction. Collectively, these results highlight that CO bond distortions and HB fluctuations actively govern the accessibility of the CoIn under acidic conditions, with the sulfate anion also playing a crucial role in this dynamic behavior.
@@
Similar CO elongation and CαCO2 deplanarization appear in the rare non-radiative trajectories of LLMHCl and LLH (Figs. S13–S14b–d), but no PT occurs in S or S. In both systems, weakening of the NHO coordinates is observed, with values often exceeding Å, but stems from other structural distortions driving the non-radiative decay. In LLMHCl, CO stretching and deplanarization lead to CαC elongation in S (Fig. S13e), distancing the acceptor from the NH donor (Fig. S13f). In LLH, significant CαN stretching near the CoIn (Fig. S14f) weakens one HB, while the other is disrupted by flipping of the carboxylate group (Figs. S14g–h). This CαN activation was recently implicated in the weak fluorescence of cysteine crystals.7 These decay events, although rare, highlight that under neutral and basic conditions, non-radiative decay is primarily driven by backbone distortions rather than HB dynamics. In contrast, under acidic conditions, the combination of enhanced protonation and the presence of flexible, multiply charged counterions–such as sulfate ions–increases the dynamical complexity of the hydrogen bonding network, opening new non-radiative pathways.
@@
In conclusion, we have combined experiments and theory to investigate the evolution of fluorescence in solid-state lysine as a function of pH. To the best of our knowledge, the physical factors and mechanisms controlling intrinsic fluorescence as a function of pH in non-aromatic supramolecular assemblies remain poorly understood and have been explored in only a handful of studies2, 32, 21. One of our key findings is that both pH and the counterions introduced by the acids used to adjust it modulate protonation states and hydrogen bonding networks, altering the physical and chemical properties of the assemblies.
@@
Our findings open up important perspectives and challenges on both experimental and theoretical fronts. Specifically, advancing experimental methods to better characterize local and macroscopic structures of titratable amino acid aggregates as a function of pH remains a crucial factor to guide computational modeling. While our models capture qualitative trends, a unique aspect we uncovered in this study is that the non-radiative decay paths under acidic conditions involve proton transfer processes for both normal and strong hydrogen bonds. An important implication of this is that the choice of the QM/MM boundary may suppress or enhance proton transfer events. While we have found that our conclusions are robust to this factor (see Fig S15 in the SI), efforts that couple reactive QM/MM boundaries for excited state dynamics will be interesting to pursue in the future.
@@
Additional experimental and computational details; determination of the critical aggregation concentration; birefringence microscopy images of Lys aggregates acidified with HCl and H2SO4 at pH 10; confocal fluorescence images of Lys aggregates prepared using HCl; normalized steady-state fluorescence spectra of Lys aggregates under varying excitation wavelengths (HCl and H2SO4); excitation-emission correlation plots; molecular orbitals involved in excitation and emission of L-Lysine sulfate (LLS), L-Lysine monohydrochloride dihydrate (LLMHCl), and L-Lysine hemihydrate (LLH); oscillator strength distributions for the three Lys crystal forms; vibrational coordinate analyses from trajectories undergoing non-radiative decay (LLS, LLMHCl, LLH); validation of the QM/MM setup for LLS; and additional references.
S1 Experimental Details
S1.1 Chemicals
L-Lysine (Lys) monohydrate, HCl (wt) and H2SO4 (wt) were purchased by Thermo Fisher Scientific (Milan, Italy).
S1.2 Samples preparation
All solutions were prepared by weight dissolving Lys in doubly distilled water; their pH was adjusted using HCl (wt) and H2SO4 (wt). pH was measured by pH-meter (Mettler Toledo FiveEasy Plus pH meter FP20, equipped with a micro electrode).
S1.3 Fluorescence spectroscopy
Fluorescence spectra were collected at room temperature on a Jasco spectrofluorophotometer (Model FP-8550), equipped with a continuous output Xenon arc lamp. Measurements in solution were conducted in a quartz cell with a 1.0 cm path length. Solid state fluorescence and photoluminescence quantum yields (PLQYs) were evaluated using an ILF-835 integrating sphere. PLQY was measured by the absolute method thereby overcoming the issue of unavoidable differences in the relative emission intensities arising from poorly reproducible drop-casted amino acid clusters. Fluorescence spectra of Lys at the solid state were acquired in a window up to 600 nm, probing the 330–410 nm excitation range at 10 nm steps. All acquisitions were averaged over five measurements after spectral correction.
S1.4 Confocal microscopy
After drop-casting on a microscope slide the Lys water solutions at various states of protonation and letting them dry overnight, samples were analyzed using a and magnifications on the MICA microhub (Leica, Wetzlar, Germany) platform. Fluorescence images were obtained using filters in the spectral regions of blue (, ), green (, ), and red (, ).
S1.5 Optical microscopy
To prepare the samples, 20 L of each Lys solution was placed onto a glass microscope slide and air-dried overnight at room temperature. The obtained samples were examined using an Optech BM80Pol microscope (Milan, Italy) for both cross polarized and bright-field light.
S1.6 Scanning Electron Microscopy
Morphological analysis of samples was carried out by field emission scanning electron microscope (ThermoFisher Scientific Phenom Pro G6). 10 L of Lys solutions were drop-casted on an aluminum stub and air-dried. A 7 nm coat of gold was sputtered at a current of 25 mA for 75 sec. The gold coated samples were then introduced into the specimen chamber and the images were acquired in high vacuum mode, at an accelerating voltage of 15 kV, through the Secondary Electron Detector (SED).
S1.7 Critical Aggregation Concentration Measuraments

### v2

@@ 3.2.1 Excited State Simulations of Lysine Crystals
Our experimental characterization shows that both pH and counterion modulate the emission of solid‑state Lys aggregates, implicating changes in protonation and aggregation as key factors. However, experimental techniques alone cannot directly resolve the molecular origins of this effect, specifically, how protonation state shapes the electronic structure and ES dynamics. To address this, we performed NAMD simulations on three Lys crystal models representing distinct protonation states, as shown in Fig. 5.
@@
The structural differences among the lysine crystals reflect protonation-dependent variations that mirror the pH effects observed in aggregate morphologies, supporting their use as proxies to study solid-state Lys photophysics. As we will see shortly, these models also correctly reproduce the trends in the emission observed as a function of pH. L-Lys sulfate (LLS, Fig. 5a) adopts a folded conformation stabilized by a dense HB network between SO and the -NH and CO2H groups of lysineCapasso_1983, KRISHNAKUMAR_2010, with HBs exclusively between Lys and sulfate and no direct Lys-Lys interactions due to electrostatic repulsion. In L-Lys monohydrochloride dihydrate (LLMHCl, Fig. 5b), all amines are protonated, the C-terminus is deprotonated, and both Lys–Lys and Lys–water HBs are observed. Cl- counterions further stabilize the structure via electrostatic interactions with both the - and -amino groupswright_1962, ramesh2006growth. Finally, when L-Lys crystallizes in its zwitterionic state, it forms a hemihydrate crystal structure (LLH, Fig. 5c). The Lys molecules orient such that the head groups–namely, the amino (-NH) and carboxyl (-CO) groups–of two monomers interact, while the side chain amino groups form HBs with water molecules, resulting in an extended HB network that propagates across the crystal layersWilliams_2015, Williams_2016. Interestingly, Arnon et al. arnon2021off showed that LLH crystals exhibit fluorescence, with its hydrogen bonding network likely playing a key role in the observed optical properties. However, their work did not include atomistic simulations to probe the mechanistic origin of the emission, which we address in the present study.
It is worth noting that, while multiple protonation states of lysine coexist at a given pH (see Fig. 1), each solid-state polymorph corresponds to a specific dominant protonation state as determined crystallographicallywright_1962, Capasso_1983, Williams_2015, Williams_2016. The protonation states present in the selected crystals (LLS, LLMHCl, and LLH) coincide with the maxima of the mole fraction distribution at pH 1, 4-7, and 10, respectively, making them natural proxies for our study. We acknowledge that this modeling strategy simplifies the complexity of real samples, where mixed protonation states and structural disorder may occur, but it allows us to isolate the effect of each dominant protonation state on fluorescence and to establish clear mechanistic connections between protonation, hydrogen-bonding networks, and excited state dynamics.
@@
For each crystal, we performed 500 fs NAMD simulations, within a QM/MM framework, on 50 statistically independent initial conditions (see Computational Details). In Fig. 5a–c, the lysine dimers rendered with glossy CPK spheres represent the QM regions used in our simulations. After 500 fs, most trajectories remained in the first ES for LLH (pH = 10) and LLMHCl (pH = 7), namely 98 and 92, whereas only 28 remained in S1 for LLS (pH = 1), as shown in Fig. 6a (orange plot). The same panel shows a qualitative comparison between these fractions and the EQE values at different pH, providing an initial validation of our computational results. The consistency between the EQE and N/N trends across pH supports the accuracy of our ES dynamics. Moreover, the severe quenching observed both in the experimental and calculated results at low pH highlights the key role of protonation in modulating ES lifetimes.
@@
Fig. 6b presents the calculated fluorescence spectra of LLS (red line), LLMHCl (green line), and LLH (cyan line) averaged over the trajectories remaining in S1. Vertical lines indicate the averaged lowest absorption (excitation) transitions: blue (LLS), magenta (LLMHCl), and yellow (LLH). For all systems, excitation wavelengths cluster near nm, exhibiting a 120 nm ( eV) blue-shift relative to experiment. This discrepancy likely arises from a combination of known TDDFT limitations, such as the tendency of CAM-B3LYP to overestimate vertical excitation energiesShao_2020, as well as neglect of nuclear quantum effects (NQEs), and structural defects or disorderMeirzadeh_2018, Kurashvili_2022. Several previous studies have shown that the inclusion of NQEs can alter both the peak absorption in organic systems by up to 1 eV, as well as affect the line-shapeFeher_2021, Sappati_2016, Pinotsi2016, Law_2017. In addition, the size of the QM subsystem can also influence spectral positions. While our simulations employed QM dimers as a compromise between accuracy and computational cost, benchmarks in other amino acid crystalsBanerjee_2025 have shown that extending to larger clusters, such as QM heptamers, red-shifts absorption energies by eV, thereby reducing the discrepancy with experiment.
@@
Across protonation states, the transition consistently occurs from the highest occupied molecular orbital (HOMO) to the lowest unoccupied molecular orbital (LUMO), which exhibit n and character, respectively. In LLS (Fig. 6c), the n-type HOMO localizes on the C-terminal carboxyl and partly on SO oxygens, while the -type LUMO resides on the same Lys monomer, primarily on the C-terminal oxygen and along the CCα bond. Similar localizations occur for LLMHCl (Fig. S7a) and LLH (Fig. S8a), with the n-type HOMO on CO oxygens and the -type LUMO on the same group and adjacent Cα. Cl- ions show no active role in LLMHCl. In the LLH crystal, both frontier orbitals are confined to the C-terminus engaged in head-to-head HBs, with no contribution from the –NH2 side chain group within our QM description. Therefore, our results indicate that the dominant photoexcited electronic structure arises from the C-terminal hydrogen bonding motif. Finally, in all systems, the electron density alternates between the two QM Lys monomers, consistent with dimer symmetry, and localizes within H-bonding regions, linking local non-covalent interactions with photoexcited electronic properties. Comparison with the isolated monomer (shown for LLH as an example in Fig. S9) highlights that the solid-state packing and HB network substantially influence the nature of the excitation. Removing the crystal environment converts the transition to an n character, with HOMO and LUMO localized on the C- and N-terminus, respectively, of the same Lys monomer, illustrating the role of aggregation in shaping the electronic structure.
@@
While excitation energies are relatively pH-independent, emission spectra show more pronounced differences. The computed emission trend in Fig. 6b mirrors the one observed experimentally for Lys samples acidified with H2SO4 (Fig. 4b). LLS (pH 1) exhibits the most blue-shifted emission ( nm), followed by LLMHCl (pH 7) and LLH (pH 10), both exhibiting emission maxima near 356 nm. LLMHCl displays an additional broad 460-480 nm feature, originating from structures in which ES distortions–particularly CCO2 deplanarization–are partially activated but insufficient to reach the conical intersection (CoIn). This mode, typically associated with non-radiative decaymiron2023, in this case gives rise to red-shifted emission from distorted yet emissive geometries. LLS emission instead exhibits a shoulder near the excitation energy, suggesting an ES minimum near the Franck-Condon region. Although emission energies remain overall blue-shifted by 50–80 nm (0.5-0.7 eV) relative to experiment, this discrepancy is significantly reduced compared to excitation, and spectral tails at longer wavelengths ( nm) are well captured for LLMHCl and LLH. All SS0 transitions retain the same n– character as the GS excitations (Figs. S7, S8, and S10).
To facilitate comparison with the normalized experimental emission spectra (see above), the calculated emission spectra in Fig. 6b were also normalized. However, normalization prevents direct comparison of absolute intensities. Therefore, we examined the distributions of oscillator strength (f) values, which measure the radiative probability of electronic transitions, also complementing the previous S1 population analysis (Fig. 6a). Normalized histograms and Gaussian fits for LLS, LLMHCl, and LLH are shown in Fig. S11a-c, with an overlay of the fitted curves in Fig. S11d. For LLH (cyan) and LLMHCl (green), the distributions peak below 0.001, with minor tails up to 0.005. Although these values are modest–well below those typical of aromatic fluorophoresGDM_2024–their overall contribution to the emission is amplified by the large fraction of trajectories remaining in S1 (Fig. 6a). In contrast, LLS (red) shows a broader distribution extending to 0.01, but its emission remains minimal because only a small fraction of trajectories survive in S1 long enough to emit, due to dominant non-radiative decay pathways.
3.2.2 Vibrational Modes Initiating Non-Radiative Decay
@@
To understand why the LLS crystal exhibits particularly weak fluorescence, we analyzed the vibrational modes that funnel the system to the CoIn and, subsequently, enable decay to the ground state. Previous studiesPinotsi2016, Sijo_2019, Grisanti2020, Stephens2021, miron2023 have identified proton transfer, carbonyl stretching, and deplanarization of the amide group as key modes governing the ES lifetimes of peptide chains, amyloid-like peptides, and glutamine crystals. Constraining these modes raises the energy barrier to the CoIn, thereby prolonging the ES lifetime. More recently, Banerjee and co-workersBanerjee_2025 linked the weaker fluorescence of cysteine crystals grown in H2O, as opposed to D2O, to activation of CαN and SH stretching modes.
Guided by the n character of LLS (Fig. 6c), we first focused on vibrational modes involving the C-terminus (Figs. S12-S13 for the two QM Lys monomers). We then analyzed coordinates associated with the system hydrogen bonding network (Fig. S14). From this combined set, a few modes emerge as the dominant funnels to the CoIn, as shown in Fig. 7.
@@
The collective motion driving non-radiative decay in LLS (Fig. 7a) can be decomposed into two main contributions: (i) stretching of the CO bond coupled with deplanarization of the CαCO2 group, and (ii) PT coordinates involving NH or OH groups of Lys and sulfate oxygens. As shown in Fig. 7b, the CO bond elongates in S1 (orange distribution), increases further at the CoIn (dashed line), and contracts upon relaxation to S0 (blue distribution). A similar trend is observed for the second CO bond of the carboxylate group (Fig. S13c). This distortion is concerted with amide deplanarization in S1, followed by replanarization after decay (Fig. S13d). These features are consistent with previously reported ES dynamics in non-emissive amino acid crystals, such as L-glutaminemiron2023.
@@
However, unlike glutamine–where CO stretching is triggered by transient strengthening (PT coordinate ) of a -NHCO HB between monomers–LLS exhibits HBs exclusively with sulfate counterions. Consequently, CO bond activation and proton transfer may occur in different structural units. The primary HB-related mode promoting non-radiative decay involves transient weakening of an -NHSO HB in S1, particularly near the CoIn, with PT typically occurring upon relaxation to S0 (Figs. 7c and S14f). A secondary PT mode involves the C-terminal OH group and displays a slight transient HB strengthening in S1, with occasional proton sharing or transfer to the sulfate oxygen at the CoIn, and reversion after decay (Fig. 7d). Many trajectories that reach the CoIn show PT coordinates below Å, reflecting a weakened interaction. Collectively, these results highlight that CO bond distortions and HB fluctuations actively govern the accessibility of the CoIn under acidic conditions, with the sulfate anion also playing a crucial role in this dynamic behavior.
@@
Similar CO elongation and CαCO2 deplanarization appear in the rare non-radiative trajectories of LLMHCl and LLH (Figs. S15–S16b–d), but no PT occurs in S or S. In both systems, weakening of the NHO coordinates is observed, with values often exceeding Å, but stems from other structural distortions driving the non-radiative decay. In LLMHCl, CO stretching and deplanarization lead to CαC elongation in S (Fig. S15e), increasing the NH–acceptor distance (Fig. S15f). In LLH, significant CαN stretching near the CoIn weakens one HB, while the other is disrupted by flipping of the carboxylate group (Figs. S16f–h). This CαN mode was recently implicated in the weak fluorescence of cysteine crystals.Banerjee_2025
These decay events, though rare, indicate that under neutral and basic conditions, non-radiative decay is primarily driven by backbone distortions rather than HB dynamics. In contrast, under acidic conditions, the combination of enhanced protonation and the presence of flexible, multiply charged counterions–such as sulfate ions–increases the dynamical complexity of the hydrogen bonding network, opening new non-radiative pathways.
4 Conclusions
@@
In conclusion, we combined experiments and theory to investigate the evolution of fluorescence in solid-state lysine as a function of pH. To the best of our knowledge, the physical factors and mechanisms controlling intrinsic fluorescence in non-aromatic supramolecular assemblies under varying pH remain poorly understood and have been explored in only a few studiesPinotsi2016, stagi2022, Wang_2025. One of our key findings is that both pH and the counterions introduced by the acids used to adjust it modulate protonation states and hydrogen bonding networks, altering the physical and chemical properties of the assemblies.
@@
Our findings open up important perspectives and challenges on both experimental and theoretical fronts. Specifically, advancing experimental methods to better characterize local and macroscopic structures of titratable amino acid aggregates as a function of pH remains a crucial factor to guide computational modeling. While our models capture qualitative trends, a unique aspect we uncovered in this study is that the non-radiative decay paths under acidic conditions involve proton transfer processes for both normal and strong hydrogen bonds. An important implication of this is that the choice of the QM/MM boundary may suppress or enhance proton transfer events. While we have found that our conclusions are robust to this factor (see Fig. S17 in the SI), efforts that couple reactive QM/MM boundaries for excited state dynamics will be interesting to pursue in the future. Moreover, in all the investigated crystals, we have limited the QM region to Lys dimers, which sufficed to capture the mechanistic trends, in line with previous studiesBanerjee_2025. Future computational work using larger QM regions or including reactive water molecules could provide improved quantitative agreement with experiments, while further elucidating the atomistic origins of intrinsic fluorescence in non-aromatic supramolecular assemblies.
@@
Critical aggregation concentration measurements; birefringence microscopy images of Lys aggregates acidified with HCl and H2SO4 at pH 10; confocal fluorescence images of Lys aggregates prepared using HCl; normalized steady-state fluorescence spectra of Lys aggregates under varying excitation wavelengths (HCl and H2SO4); excitation-emission correlation plots; molecular orbitals involved in excitation and emission of L-Lysine sulfate (LLS), L-Lysine monohydrochloride dihydrate (LLMHCl), and L-Lysine hemihydrate (LLH); oscillator strength distributions for the three Lys crystal forms; vibrational coordinate analyses from trajectories undergoing non-radiative decay (LLS, LLMHCl, LLH); validation of the QM/MM setup for LLS; and additional references.
S1 Supplementary Text
S1.1 Critical Aggregation Concentration
S1.1.1 Experimental Details